# Time Series Foundation Models: Benchmarking Challenges and Requirements

### v1

@@ 5 TSFM Benchmark Requirements
The essence of time series forecasting is to predict the future. Paradoxically, using data from the past to evaluate their ability to do so is standard practice as we have seen in previous Section 4. Before the advent of TSFMs, this paradox was not a serious issue, as traditional time series models have to be trained on past values of the same time series they are expected to continue. The only pitfall to avoid is adhering to the correct ordering of training and test observations. Yet, as we have outlined in Section 2, the possibility of pre-training TSFMs on vast collections of time series from different domains and using them without training on the target time series introduces new risks of information leakage, possibly resulting in test set contamination and memorization of global patterns. Now there are a variety of ways (Section 4) to test and compare TSFM, but in our opinion none of them fully solves the problems identified (Section 2).
Instead of only pointing out which benchmarking methodologies are not suitable for TSFMs, we want to be proactive and develop guidelines how a TSFM benchmark should be designed. Based on the advantages and disadvantages of the current benchmarking approaches (Section 4) combined with our analysis (Section 2), we propose the following requirements for a time series foundation model benchmark methodology:
  1. 1.
    Prevention of Information Leakage. The benchmark must guarantee that no portion of test data is present within any model’s training corpus.
  2. 2.
    Unambiguous Data Splits and Transparency. It should employ clear and standardized protocols for dividing data into training and test sets.
  3. 3.
    Time- and Domain Dimension Evaluation. It should apply cross-validation across both time and domain dimensions, verifying model consistency across seasonal periods (e.g., summer/winter) and across diverse datasets or domains to ensure generalization and robustness.
  4. 4.
    Retrospective and Prospective Evaluation. The methodology should not be limited to models developed after the introduction of the benchmark, but must also be applicable to existing models and future models as the field evolves.
  5. 5.
    Flexibility in Pre-Training Data. Researchers should not face artificial restrictions on the amount, provenance, or specificity of pre-training data.
  6. 6.
    Mitigation of Global Pattern Memorization. Mechanisms must be in place such as global temporal cut-off points to reduce the risk of models exploiting global patterns.
  7. 7.
    Scalability and Sustainability. The framework should remain practical for long-term community use, avoiding excessive computational, infrastructural, or administrative requirements.
  8. 8.
    Real-World Representativeness. Evaluation data should reflect the diversity of time series in relevant real-world domains and frequencies.
  9. 9.
    Support for Zero-Shot and Fine-Tuned Evaluation. The protocol should enable assessment of models in both zero-shot and fine-tuned settings.
  10. 10.
    Inclusivity and Accessibility. Participation should be open to both academic and industry teams, including those using open-source or closed-source models.
  11. 11.
    Robustness Against Gaming and Overfitting. Safeguards must be implemented to prevent exploitation of the test data, such as rotating test series, or anonymizing series during evaluation.
Now, one could argue that these are wishful requirements for a benchmark, but that they are unrealistic to fulfill all and compromises must be made. We would like to challenge precisely this and encourage the community to look for solutions that cover as wide a range of requirements as possible.
For example, a benchmark based entirely on constantly generated synthetic data could meet more than half of the 11 requirements. While there has been impressive progress in mimicking real-world data (Narasimhan et al., 2024), a question arises: as the generation process becomes more similar to actual historical data, does the risk of information leakage or memorizing global patterns increase? On the other hand, if the generation process is more independent of past data, we once again encounter the issue of lacking real-world relevance. Additionally, such synthetic data may be vulnerable to gaming if someone can reverse-engineer the data generation process. Nevertheless, it is worth considering this approach further.
To more comprehensively address the spectrum of requirements, a conceptual shift in how TSFM benchmarks are designed might be necessary. Imagine a framework built around a continuously advancing global temporal split. In such a system, TSFMs would be tasked with forecasting actual future events, with evaluations performed retrospectively as real-world data becomes available. This dynamic approach, where the “test set" is always the yet-unseen future, inherently addresses critical issues like information leakage from test set contamination or rote memorization. This evaluation seems especially applicable for time-series forecasting, which is uniquely concerned with future prediction unlike typical NLP or computer vision tasks.
Furthermore, since the test data would not exist during model training, restrictions on training data could be completely lifted, facilitating the evaluation of a broader range of models, including closed-source TSFMs. Still, this paradigm introduces new complexities. For instance, distinguishing purely zero-shot forecasts from fine-tuned forecasts could be difficult and leading to nonfulfillment of requirement 9. Here it might require complementary ideas. To further promote transparency regarding training data, the adoption of model cards could be encouraged as a best practice, though community acceptance, as seen with LLMs, might present a hurdle.
Undoubtedly, operationalizing such a benchmark would present significant implementation challenges, likely necessitating a dedicated platform. However, we posit that pursuing this line of thinking offers a promising avenue to resolve many of the current impasses in TSFM evaluation. We therefore call upon the research community to collaboratively explore and develop these concepts further, shaping a more robust and reliable future for time series forecasting benchmarks.
@@ 6 Conclusion
Remarkably, TSFMs appear to follow a development trajectory similar to that of LLMs in terms of both data and model size. Scaling laws seem to apply to TSFMs, and there are notable similarities in the evaluation challenges they face, particularly concerning potential information leakage. We highlight key evaluation challenges that arise from the emergence of TSFMs. By investigating the data lineage of 15 prominent TSFMs and reflecting the common TSFM training and evaluation strategy, we identified two significant sources of information leakage: test set contamination through multi-purpose use of datasets and memorization of global patterns induced by external events or crises. Furthermore, we question the current evaluation strategy that considers only the domain dimension while disregarding the temporal dimension, as well as the selection of time series used to represent a domain.
Current benchmark approaches from time-series or machine learning do not mitigate all information leakage risks. Based on our investigations and existing methodologies, we propose requirements for a benchmark in the age of TSFMs. We would like to encourage the community to develop new ideas which can fulfill the requirements like benchmarking on real future data. Besides the development of a benchmark methodology, we see several directions for future research that could also reinterpret our current results. The effect size of test set contamination (when a test dataset was present in the pretraining corpus) has only been proven to be significant in isolated cases so far. The magnitude of global pattern memorization is also largely unknown and is likely to vary depending on the domain, type of global event, and other factors. Carefully designed training regimes for TSFMs will be required to properly quantify these effects.
Ultimately, we believe that the criteria we have established lay the foundation for robust, fair, and information-leakage-free benchmarking. This is essential for preventing an evaluation crisis for TSFMs akin to what has been observed with LLMs.
@@ Appendix A Time Series Foundation Model Analysis
A.1 Overlap in test data usage
Table 2: Comparison of train/test evaluations of selected different models ordered by the total number of overlaps. For full results see Section C.
@@ A.1 Overlap in test data usage
Table 3: Comparison of zero-shot evaluations of selected different models ordered by the total number of overlaps. Only datasets with 5 or more overlaps are shown. Models which have not included any of the shown datasets into their train/test evaluation are not displayed. For full results see Section C.
@@
A.2 Data Analysis
Additional visualizations based on the analysis of Time Series Foundation Models training and evaluation data.
Figure 5: Most common frequencies in investigated datasets.
@@ Appendix B Leakage Investigations
B.1 Potential global pattern leakage in dataset collections
It may be theoretically possible that existing model evaluation results have already been influenced by global pattern leakage. We analyse the overlapping time frames in the Monash Dataset  [^6] https://huggingface.co/datasets/Monash-University/monash_tsf as can be seen in Figure 7. Test datasets shouldn’t lie in the same time period as other training datasets. As an example, the proposed Australian Electricity Demand test time frame (April 2015) lays in the same period as the training period of the "Weather (Australia)" training set (1900-2021). The Moirai and Moirai-MoE models use both datasets for train/test forecasts, meaning if their are any global patterns in the weather data which could influence the Australian Electricity Demand data, this would be a kind of information leakage. Similarly, the Oikolab Weather dataset ranges from 2010 January to 2021 May, which is weather data located in Melbourne. The test set of the dataset Pedestrian Count (also Melbourne) is located at the end of April 2020. This test period lays within the training data of the Oikolab Weather, which could be a greater potential global pattern information leakage than in previous example. Both datasets are included as well in the train/test datasets of Moirai and Moirai-MoE.
Moreover, LagLlama uses the Beijing Multi-Site Air Quality dataset from 12 districts in Beijing from 2013 to 2017 as a pre-training dataset and simultaneously uses the Beijing PM2.5 air quality dataset from the Beijing airport between 2010 and 2013.
B.2 Leakage in Moirai Model
@@
The investigated data includes 9 datasets on short horizons, four datasets each on medium (10x short horizon) and long horizons (15x short horizon). We filtered the results to 4 datasets where data is available for all horizons to reasonably compare the metrics and model sizes, as shown in Figure 8. On short horizons, both models perform better with increasing model size while Moirai Leakage has on average 7.6% points lower MAPE score. At medium horizons, the leakage-free MoiraiL generally underperforms compared to other model sizes, exhibiting an average performance deficit of 32% points compared to the Moirai LeakageL model. On long horizons, all models perform significantly worse. However, Moirai Leakage achieves an average MAPE score that is 29% points lower than the leakage-free Moirai. Across all horizons, the biggest difference is between MoiraiL and Moirai LeakageL, similar to typical LLM memorization behavior, when large models better memorize pre-train data than smaller models.
@@
Appendix C TSFM Data Lineage
Table 5: Analysis of datasets used by investigated TSFMs with Pre-training (P), train-test splits (T/T) and test (T).
@@ Appendix C TSFM Data Lineage
Dataset Source Domain Frequency Chronos TimesFM Moirai Moirai-MoE Time-GPT ForecastPFN MOMENT Lag-Llama GPT4TS Timer ZSisionTS TTM LLMTime Time-LMM UniTime Time-MoE
Jena Weather Autoformer (Wu et al., 2021) Nature 10T ZS ZS T/T T/T ZS ZS ZS ZS T/T
Jena Weather Autoformer (Wu et al., 2021) Nature H
Jena Weather Autoformer (Wu et al., 2021) Nature D ZS
BizITObs - Application AutoMixer (Palaskar et al., 2024) Web/CloudOps 10S ZS
BizITObs - L2C AutoMixer (Palaskar et al., 2024) Web/CloudOps 5T ZS
BizITObs - L2C AutoMixer (Palaskar et al., 2024) Web/CloudOps H
BizITObs - Service AutoMixer (Palaskar et al., 2024) Web/CloudOps 10S
BDG-2 Bear BuildingsBench (Emami et al., 2023) Energy H P P P
BDG-2 Fox BuildingsBench (Emami et al., 2023) Energy H P P P
BDG-2 Panther BuildingsBench (Emami et al., 2023) Energy H P P P
BDG-2 Rat BuildingsBench (Emami et al., 2023) Energy H P P P
Borealis BuildingsBench (Emami et al., 2023) Energy H P P P
Buildings900K BuildingsBench (Emami et al., 2023) Energy H P P P
IDEAL BuildingsBench (Emami et al., 2023) Energy H P P P
Low Carbon London BuildingsBench (Emami et al., 2023) Energy H P P
Sceaux BuildingsBench (Emami et al., 2023) Energy H P P P
SMART BuildingsBench (Emami et al., 2023) Energy H P P P
CMIP6 ClimateLearn (Nguyen et al., 2023) Climate 6H P P P
ERA5 ClimateLearn (Nguyen et al., 2023) Climate H P P P P
Alibaba Cluster Trace 2018 CloudOpsTSF (Woo et al., 2023) CloudOps 5T P P P
Azure vM Traces 2017 CloudOpsTSF (Woo et al., 2023) CloudOps 5T P P P
Borg Cluster Data 2011 CloudOpsTSF (Woo et al., 2023) CloudOps 5T P P P
@@
Bitbrains Fast Storage Grid Workloads Archive (Shen et al., 2015) Web/CloudOps 5T
Bitbrains Fast Storage Grid Workloads Archive (Shen et al., 2015) Web/CloudOps H
Bitbrains Fast Storage Grid Workloads Archive (Shen et al., 2015) Web/CloudOps 5T
Bitbrains Fast Storage Grid Workloads Archive (Shen et al., 2015) Web/CloudOps H
@@
M5 GluonTS (Alexandrov et al., 2020a) Sales D ZS P P P
Taxi GluonTS (Alexandrov et al., 2020a) Transport 30T T/T P P P
Taxi GluonTS (Alexandrov et al., 2020a) Transport H P P
Uber TLC Daily GluonTS (Alexandrov et al., 2020a) Transport D T/T P P P
Uber TLC Hourly GluonTS (Alexandrov et al., 2020a) Transport H T/T P P P P
Wiki-Rolling GluonTS (Alexandrov et al., 2020a) Web D P
Trends hourly Google Trends Web H P
Trends weekly Google Trends Web W P
Trends monthly Google Trends Web D P
@@
ETT1 Informer (Zhou et al., 2020) Energy 15T ZS ZS T/T P T/T ZS ZS ZS T/T T/T T/T
ETT1 Informer (Zhou et al., 2020) Energy H ZS ZS T/T P T/T ZS ZS ZS T/T T/T T/T
ETT1 Informer (Zhou et al., 2020) Energy D ZS ZS
ETT1 Informer (Zhou et al., 2020) Energy W-THU
ETT2 Informer (Zhou et al., 2020) Energy 15T ZS ZS T/T ZS T/T ZS ZS ZS T/T T/T T/T
ETT2 Informer (Zhou et al., 2020) Energy H ZS ZS T/T P T/T ZS ZS ZS T/T T/T T/T
ETT2 Informer (Zhou et al., 2020) Energy D ZS ZS
ETT2 Informer (Zhou et al., 2020) Energy W-THU
Weather (US) Informer (Zhou et al., 2020) Weather H ZS
Brazilian Cities Temperature Kaggle Competition Weather M P

### v2

@@ 4 TSFM Benchmark Requirements
The essence of time series forecasting is to predict the future. Paradoxically, using data from the past to evaluate their ability to do so is standard practice (Section 2.2). Before the advent of TSFMs, this paradox was not a serious issue, as traditional time series models have to be trained on past values of the same time series they are expected to continue. The only pitfall to avoid is adhering to the correct ordering of training and test observations.
However, the paradigm shift behind TSFMs introduces new possibilities for information leakage. Unlike local models, TSFMs learn patterns globally. Consequently, strict temporal ordering within a single time series is no longer sufficient to guarantee a fair evaluation. To rigorously assess these models, we must address leakage across two distinct dimensions: the sample dimension and the temporal dimension (see Figure 5).
Figure 5: Information leakage risks and challenges arising from the advent of TSFMs, and the requirements needed to address them.
Flexibility in Pre-Training Data (Training Dimension) The current lack of transparency surrounding TSFM pre-training corpora creates a substantial risk of information leakage as shown in Section 3.1. This growing confusion is evident in newer benchmarks and models that report results contaminated by data leakage—whether unintentional or intentional Aksu et al. (2024); Auer et al. (2025); Li et al. (2024b). Attempting to mitigate this risk by restricting pre-training data is both counterproductive and ineffective: such restrictions artificially limit the potential of foundation models and, given the opacity of large-scale pre-training corpora, our lineage analysis shows that it is almost impossible to rule out train-test sample overlaps. We therefore argue for full flexibility in pre-training data. A robust evaluation framework should be agnostic to pre-training data choices, allowing the use of any existing data to fully exploit the predictive capabilities of TSFMs. This perspective shifts the responsibility for preventing leakage from the TSFM developer to the benchmark designer, thereby necessitating evaluation datasets that are genuinely novel and capable of assessing generalization rather than memorization.
Strict Global Post-Training Test Periods (Inference Dimension) Indirect temporal dependencies between time series, originating from direct or indirect causal relationships, create another more subtle risk for information leakage. Preventing this risk by enforcing global cut-off dates for pre-training data to separate training from test data is not a viable solution to this challenge. First, it would require all researchers to tailor their training protocol to exactly this date and second, it prevents new time series patterns of current time periods after the cut-off from being trained into the models. We therefore argue for strict global post-training test periods. These test periods need to be flexibly determined based on the global last pre-training time point of all evaluated models, instead of forcing a pre-training cut-off. For a robust evaluation framework a strict temporal barrier must be enforced across the entire universe of data: if the last pre-training timestamp across all evaluated models is $t$, only observations from $t+1$ onwards can be present in the test set of a benchmark, ensuring that TSFMs rely solely on past dynamics.
Addressing these two dimensions (unrestricted training data and strict temporal inference separation) is non-trivial and often overlooked in current evaluation standards. We ask the community to develop ideas for new benchmark methodologies which are grounded in these two requirements. Possible directions could be constantly newly sourced test data, synthetic test data, or fast-paced competitions on unreleased and unbiased test data, just to name a few.
@@ 5 Conclusion
Remarkably, TSFMs appear to follow a development trajectory similar to that of LLMs in terms of both data and model size. Scaling laws seem to apply to TSFMs, and there are notable similarities in the evaluation challenges they face, particularly concerning potential information leakage. We highlight key evaluation challenges that arise from the emergence of TSFMs. By investigating the data lineage of 22 prominent TSFMs and reflecting on common TSFM training and evaluation strategies, we identified two significant sources of information leakage: direct information leakage by train-test sample overlaps through multi-purpose use of datasets and indirect information leakage by temporal overlap of correlated train and test series.
Current benchmark approaches from time series forecasting or machine learning do not mitigate all information leakage risks. Based on our investigations and existing methodologies, we propose two key requirements for information leakage free evaluations in the age of TSFMs: flexibility in pre-training data and strict global post-training test periods. Beyond the development of new benchmark methodologies, we identify several directions for future research . To date, the effect size of test-set contamination—when a test dataset is present in the pre-training corpus—has only been shown to be significant in isolated cases. Likewise, the magnitude of temporal leakage arising from correlated and overlapping time series remains largely unknown in real TSFM settings and is likely to vary across domains, types of global events, and other contextual factors. Carefully designed TSFM training regimes will be required to rigorously quantify these effects.
Ultimately, we believe that the requirements we have established lay the foundation for robust, fair, and information-leakage-free benchmarking. This is essential for preventing an evaluation crisis for TSFMs akin to what has been observed with LLMs.
6 Data Availability
The resulting table of the TSFM data-lineage analysis is supplied in the Supplementary Information Section D. For the global pattern experiment in Section 3.2, the used data is available through the source code via zenodo: https://doi.org/10.5281/zenodo.18635478
7 Code Availability
The code to reproduce the experiment in Section 3.2 is provided via zenodo: https://doi.org/10.5281/zenodo.18635478
@@ Appendix A Time Series Foundation Model Analysis
A.1 Multi-Purpose Use of Train and Test Data
This section showcases the varying training and testing dataset strategies of selected TSFMs. Table 3 reveals that there is virtually no standardized approach among TSFMs, with each model employing different combinations of datasets for train/test (or in-domain) datasets. While some datasets like Electricity (H) and Traffic Hourly (H) appear across multiple models, the coverage remains fragmented. Table 4 further illustrates this inconsistency in zero-shot evaluation practices. Even among datasets with five or more overlaps, the specific models testing on each dataset vary considerably. For instance, ETT2 (15T) is evaluated by Chronos, Moirai, and Lag-Llama, while Electricity (H) appears in Moirai, Moirai-MoE, and LLMTime.
Another notable case for the multi-purpose use of datasets is the subset Australian Electricity Demand from the Monash dataset collection (Godahewa et al., 2021), which is used for pre-training (e.g., Lag-Llama, Timer), train/test evaluation (e.g., Moirai), and zero-shot forecasting (e.g., Chronos, TimesFM). This multi-purpose use makes the dataset practically useless for benchmarking between models (e.g., it cannot be used to compare Chronos to Lag-Llama). To understand the problem of overlaps better, we provide Figure 6.
Also an interesting case is the Informer collection (Zhou et al., 2021), which is primarily used for zero-shot evaluations. However, some individual time series of the Informer collection, such as ETTh1, ETTh2, and ETTm1, have also been used for pre-training (e.g., in LagLlama) or train/test evaluation (e.g., in UniTime). This means that LagLlama or UniTime cannot be benchmarked against any other model using the Informer collection.
Taken together with investigations of Roque et al. (2025) that the selection of evaluation datasets can highly influence the final evaluation results, this strengthens the need for a more uniform benchmark approach.
Table 3: Comparison of train/test evaluations of selected different models ordered by the total number of overlaps. For full results see Section D.
@@ A.1 Multi-Purpose Use of Train and Test Data
Table 4: Comparison of zero-shot evaluations of selected different models ordered by the total number of overlaps. Only datasets with 5 or more overlaps are shown. Models which have not included any of the shown datasets into their train/test evaluation are not displayed. For full results see Section D.
@@ Appendix B Leakage Investigations
B.1 Effects of Overlaps in Train-Test Samples
This section collects different examples where the train-test sample overlaps unintentionally led to information leakage in recent benchmark papers. In one case, the benchmark creators, due to the TSFMs’ incomprehensible datasets, unintentionally included three evaluation datasets as test sets that had already been used in the pretraining of TimesFM, UniTS, and TTM (Li et al., 2024b). Based on our analysis of the benchmark results, this leads to an advantage of 47% - 184% lower mean squared error (MSE) rate compared to best models not pre-trained on the leaking datasets. The advantage of the best TSFM on non-leaked datasets is only between 0.3% and 14%. The performance benefit through information leakage appears to align with findings from the Moirai leakage example (see Section B.2). In another example a peer-reviewed paper benchmarked TSFM models on the Electricity dataset which is included in every tested TSFM pretraining data Saravanan et al. (2024). Also in the energy domain, a paper benchmarked TSFMs on the Spanish dataset (among others), which is also included in the pretraining data of the tested TSFMs. Chronos achieved a sMAPE of 4.854 where the best non-TSFM model TiDE achieved a sMAPE of 8.102 Montet et al. (2025). This highlights the increasing difficulties for benchmarking TSFMs.
B.2 Intended Information-Leakage Experiment for Moirai Model
@@
The empirical analysis using the Moirai TSFM by Aksu et al. (2024) demonstrates the effects of information leakage between training and test sets on the predictive performance of TSFMs. The authors prepared a new pre-training and held-out evaluation dataset without any overlap and trained the Moirai architecture on it. A different Moirai model from a previous publication was taken which pre-training dataset contained 0.1% data of the newly defined held-out evaluation dataset leading to a deliberate small portion of information-leakage. The investigated data includes nine datasets tested on three different forecast horizons (short, medium, long). Not for all horizons data was available, so we filtered the results on 4 datasets where the results are available for all horizons to reasonably compare the metrics and model sizes, as shown in Figure 7. On short horizons, both models, leakage-free and leakage model, perform better with increasing model size while Moirai Leakage has on average 7.6% points lower MAPE score. At medium horizons, the leakage-free MoiraiL generally underperforms compared to other model sizes, exhibiting an average performance deficit of 32% points compared to the Moirai LeakageL model. On long horizons, all models perform significantly worse. However, Moirai Leakage achieves an average MAPE score that is 29% points lower than the leakage-free Moirai. Across all horizons, the biggest difference is between MoiraiL and Moirai LeakageL.
These results vividly demonstrate the significant impact that train-test sample overlap has on the performance of TSFMs. Larger model sizes especially seem prone to this.
@@
B.3 Potential Overlaps of Correlated Train and Test Series in Dataset Collections
It may be theoretically possible that existing model evaluation results have already been influenced by overlaps of correlated train and test series. We analyse the overlapping time frames in the Monash Dataset  [^10] https://huggingface.co/datasets/Monash-University/monash_tsf as can be seen in Figure 8. Test datasets shouldn’t lie in the same time period as other training datasets. As an example, the proposed Australian Electricity Demand test time frame (April 2015) lays in the same period as the training period of the "Weather (Australia)" training set (1900-2021). The Moirai and Moirai-MoE models use both datasets for train/test forecasts, meaning if their are any global patterns in the weather data which could influence the Australian Electricity Demand data, this would be a kind of information leakage, at least on theoretical basis. Similarly, the Oikolab Weather dataset ranges from 2010 January to 2021 May, which is weather data located in Melbourne. The test set of the dataset Pedestrian Count (also Melbourne) is located at the end of April 2020. This test period lays within the training data of the Oikolab Weather, which could be a greater potential global pattern information leakage than in previous example. Both datasets are included in the datasets of Moirai, Time-Moe, Sundial and Toto.
Moreover, LagLlama uses the Beijing Multi-Site Air Quality dataset from 12 districts in Beijing from 2013 to 2017 as a pre-training dataset and simultaneously uses the Beijing PM2.5 air quality dataset from the Beijing airport between 2010 and 2013, which could also in theory lead to leakage.
Appendix C Temporal Overlap of Correlated Series Experiment
This section provides implementation details of the experiment conducted to show the influence of the temporal overlap of correlated train and test series in a TSFM setting.
C.1 Data visualizations and statistics
Figure 9 shows two example time series from major stock indices, the DAX and S&P 500. Despite representing different markets, both series display remarkably similar patterns during key economic phases, including crisis periods and growth cycles. This visual comparison highlights how global financial markets often move in tandem, reflecting shared responses to worldwide economic events.
Figure 10 illustrates the relationship between training data and test data that overlap temporally. It is important to note that while the training and test periods overlap in time, they do not share the same data points, the overlap is purely temporal, meaning both datasets cover in a small part the same time period but contain different observations or series as can be seen in Table 6.
Figure 9: Comparison of DAX and S&P 500 points over time
Figure 10: Comparison of time periods per dataset
Category No Leakage Indirect Leakage
Train Test Train Test
# Time Series 484 1 484 1
# Data Points 205,854 230 205,859 230
# Indirect Leakage Data Points - - 1,600 -
Table 6: Summary of time series and data points with and without leakage
C.2 Scaling and Model Tuning
Min-max scaling was applied separately to each dataset to account for differences in scale across datasets and ensure optimal model performance. Each scaler was fitted on the respective training data and subsequently used to transform the remaining data. For the S&P 500 dataset, the scaler was fitted explicitly on data from 2015 to the end of 2019 and then applied to the data from 2020. The Time-Series Transformer was implemented via the tsai library [^11] https://timeseriesai.github.io/tsai. The tuning was performed via Optuna [^12] https://github.com/optuna/optuna for 50 trials to minimize the MAE on the validation set. The training set consisted of the first 80% of NN5 Daily and Tourism Monthly datasets as well as January 1st, 2015 until June 30th, 2019 for no-leak and January 1st 2016 until June 30th, 2020 of the stock index data for leak. The validation set consisted of the remaining 20% of NN5 Daily and Tourism Monthly datasets, and the stock index data from July 1st, 2019 until December 31st, 2019 for no-leak and July 1st, 2020 until December 31st 2020 for leak.
Parameter Lower Upper Stepsize Choice
Learning rate log(1e-05) log(1e-01) - -
Batch size - - - 16, 32, 64, 128
Epochs 5 30 5 -
Dropout 0.0 0.7 - -
FC Dropout 0.0 0.7 - -
Number of layers 2 8 1 -
Model dimension - - - 64, 128, 256, 512, 1024
Number of heads - - - 4, 8, 16, 32
Dimension of feedforward network model - - - 32, 64, 128, 256, 512
Table 7: Optuna search space for tuning
Parameter No-leak Leak
Learning rate 1.018e-03 1.431e-03
Batch size 128 128
Epochs 30 30
Dropout 0.5 0.3
FC Dropout 0.27 0.004
Number of layers 2 4
Model dimension 128 128
Number of heads 32 4
Dimension of feedforward network model 32 512
Table 8: Final model parameters
Appendix D TSFM Data Lineage
@@
Dataset Source Domain Frequency Chronos TimesFM Moirai Moirai-MoE Time-GPT ForecastPFN MOMENT Lag-Llama GPT4TS Timer VisionTS TTM LLMTime Time-LMM UniTime Time-MoE Chronos-2 Flowstate(under review) LightGTS Sundial TabPFN-TS Time-MoE2 TinyTimeMixer TiRex Toto YingLong (under review) Gift-Eval TOTO2 Flowstate(T) TimesFM-ICL ROSE
Jena Weather Autoformer (Wu et al., 2021) Nature 10T ZS ZS T/T T/T ZS ZS ZS ZS T/T ZS T/T ZS T/T ZS T/T
Jena Weather Autoformer (Wu et al., 2021) Nature H ZS T/T ZS
Jena Weather Autoformer (Wu et al., 2021) Nature D ZS ZS T/T ZS
BizITObs - Application AutoMixer (Palaskar et al., 2024) Web/CloudOps 10S ZS ZS T/T ZS
BizITObs - L2C AutoMixer (Palaskar et al., 2024) Web/CloudOps 5T ZS ZS T/T ZS
BizITObs - L2C AutoMixer (Palaskar et al., 2024) Web/CloudOps H ZS T/T ZS
BizITObs - Service AutoMixer (Palaskar et al., 2024) Web/CloudOps 10S ZS T/T ZS
BDG-2 Bear BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
BDG-2 Fox BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
BDG-2 Panther BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
BDG-2 Rat BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
Borealis BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
Buildings900K BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
IDEAL BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
Low Carbon London BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P
Sceaux BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
SMART BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
CMIP6 ClimateLearn (Nguyen et al., 2023) Climate 6H P P P P P P P P
ERA5 ClimateLearn (Nguyen et al., 2023) Climate H P P P P P P P P P
Alibaba Cluster Trace 2018 CloudOpsTSF (Woo et al., 2023) CloudOps 5T P P P P P P P
Azure vM Traces 2017 CloudOpsTSF (Woo et al., 2023) CloudOps 5T P P P P P P P P P P
Borg Cluster Data 2011 CloudOpsTSF (Woo et al., 2023) CloudOps 5T P P P P P P P P P P
@@
Bitbrains Fast Storage Grid Workloads Archive (Shen et al., 2015) Web/CloudOps 5T ZS T/T ZS
Bitbrains Fast Storage Grid Workloads Archive (Shen et al., 2015) Web/CloudOps H ZS T/T ZS
Bitbrains Fast Storage Grid Workloads Archive (Shen et al., 2015) Web/CloudOps 5T ZS T/T ZS
Bitbrains Fast Storage Grid Workloads Archive (Shen et al., 2015) Web/CloudOps H ZS T/T ZS
@@
M5 GluonTS (Alexandrov et al., 2020a) Sales D ZS P P P P P P P ZS ZS
Taxi GluonTS (Alexandrov et al., 2020a) Transport 30T T/T P P P P P T/T P P P P
Taxi GluonTS (Alexandrov et al., 2020a) Transport H P P possible P P P P P P
Uber TLC Daily GluonTS (Alexandrov et al., 2020a) Transport D T/T P P P P P T/T P P P
Uber TLC Hourly GluonTS (Alexandrov et al., 2020a) Transport H T/T P P P P P P T/T P P P
Wiki-Rolling GluonTS (Alexandrov et al., 2020a) Web D P P P P
Trends hourly Google Trends Web H P P
Trends weekly Google Trends Web W P P
Trends monthly Google Trends Web D P P
@@
ETT1 Informer (Zhou et al., 2020) Energy 15T ZS ZS T/T P T/T ZS ZS ZS T/T T/T T/T ZS T/T ZS T/T ZS ZS ZS T/T
ETT1 Informer (Zhou et al., 2020) Energy H ZS ZS T/T P T/T ZS ZS ZS T/T T/T T/T ZS T/T ZS T/T ZS ZS ZS T/T
ETT1 Informer (Zhou et al., 2020) Energy D ZS ZS ZS T/T ZS
ETT1 Informer (Zhou et al., 2020) Energy W-THU ZS T/T ZS
ETT2 Informer (Zhou et al., 2020) Energy 15T ZS ZS T/T ZS T/T ZS ZS ZS T/T T/T T/T ZS T/T ZS T/T ZS ZS T/T
ETT2 Informer (Zhou et al., 2020) Energy H ZS ZS T/T P T/T ZS ZS ZS T/T T/T T/T ZS T/T ZS T/T ZS ZS T/T
ETT2 Informer (Zhou et al., 2020) Energy D ZS ZS ZS T/T ZS
ETT2 Informer (Zhou et al., 2020) Energy W-THU ZS T/T ZS
Weather (US) Informer (Zhou et al., 2020) Weather H ZS T/T
Brazilian Cities Temperature Kaggle Competition Weather M P possible P P P P P